Ammonium nitrite, NH4NO2, decomposes upon heating to form N2 gas according to the following balanced chemical equation. When a sample of NH4NO2 was decomposed in a test tube, 813.9 mL of N2 gas was collected over water at 47.64 °C and the total pressure was 765.9 torr. NH4NO2(s) → N2(g) + 2 H2O(l).

How many grams of N2 were collected?
Answer #1


Prone = PLATER UApor + PNG GIUEN Prome - 76518 TORR PWATER VAPOR = 8015 TORK AT 47064°c Pro = 76508 - 80.5= 685-4 TORR = 685.